# ReCalKV: Low-Rank KV Cache Compression via Head Reordering and Offline Calibration

### v2

@@ 3.3 Offline Calibration and Matrix Fusion
$\mathcal{E}=||\mathbf{L}_{v}\mathbf{R}_{v}\mathbf{X}-\mathbf{W}_{v}\mathbf{X}||_{F}^{2},$ (6)
where $\mathbf{X}$ denotes the calibration dataset. Based on our analysis of Fisher Information, we observe that the Value projection matrix exhibits significantly higher Fisher Information compared to the Key projection matrix. This indicates that the Value projection matrix plays a more critical role in model performance. Therefore, we aim to minimize the approximation error $\mathcal{E}$ introduced during the compression of the Value projection to preserve model accuracy as much as possible. Inspired by recent findings in adaptive decomposition methods such as AdaSVD li2025adasvd , we observe that standard SVD decomposition does not always yield the lowest approximation error. To this end, we perform offline calibration of the decomposed matrices $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ using a small calibration dataset $\mathbf{X}$, aiming to further reduce the compression-induced approximation error. We first calibrate $\mathbf{L}_{v}$ by setting the derivative of the approximation error $\mathcal{E}$ with respect to $\mathbf{L}_{v}$ to zero:
@@
This yields an optimal closed-form solution for $\mathbf{R}_{v}$. Together, the calibrated $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ form a refined low-rank approximation of the original Value projection matrix, effectively reducing compression-induced error without incurring any additional inference-time computation. The derivation of the calibration objective and the resulting closed-form updates for both $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ are provided in the supplementary material, along with a detailed discussion of its implementation considerations.
@@
4:     $\mathcal{F}\leftarrow$ Calculate_Fisher_Info$(\mathcal{M},\mathcal{X})$
5:     $R\leftarrow$ Allocate_Compression_Ratio$(\mathcal{M},\mathcal{TR},\mathcal{F})$
@@
7:         $\mathcal{CKA}\leftarrow$ Calculate_CKA_Similarity$(\mathcal{W}_{k})$
8:         $\mathcal{W}_{k}^{\prime}\leftarrow$ Head_Reorder$(\mathcal{W}_{k},\mathcal{CKA})$
9:         $\mathbf{L}_{k}[],\mathbf{R}_{k}[]\leftarrow$ Group_SVD$(\mathcal{W}_{k}^{\prime},\mathcal{R}[\mathcal{W}_{k}])$
10:         $\mathcal{M}^{\prime}\leftarrow$Update_Layer$(\mathbf{L}_{k}[]\&\mathbf{R}_{k}[],\mathcal{W}_{k},\mathcal{M})$
@@
13:         $\mathbf{L}_{v},\mathbf{R}_{v}\leftarrow$ SVD$(\mathcal{W}_{v},\mathcal{R}[\mathcal{W}_{v}])$
14:         $\mathbf{L}_{v}^{\prime},\mathbf{R}_{v}^{\prime}\leftarrow$ Offline_Calibration$(\mathcal{W}_{v},\mathbf{L}_{v},\mathbf{R}_{v},\mathcal{X})$
15:         $\widetilde{\mathcal{W}}_{o}\leftarrow$Matrix_Fusion$(\mathbf{R}_{v}^{\prime},\mathcal{W}_{o})$
16:         $\mathcal{M}^{\prime}\leftarrow$Update_Layer$(\mathbf{L}_{v}^{\prime},\mathcal{W}_{v},\mathcal{M})$
17:         $\mathcal{M}^{\prime}\leftarrow$Update_Layer$(\mathcal{W}_{o}^{\prime},\mathcal{W}_{o},\mathcal{M})$
@@ 3.4 ReCalKV Workflow
The overall pipeline of ReCalKV is outlined in Algorithm 1. Given a pre-trained model $\mathcal{M}$, calibration data $\mathcal{X}$, and a target compression ratio $\mathcal{R}$, ReCalKV applies differentiated compression strategies to Key and Value projection layers. We first compute layer-wise Fisher Information scores using the calibration data, following the strategy introduced in Palu chang2024palu , to estimate the relative importance of each layer and guide compression ratio allocation. Based on this, we allocate compression ratios to different layers accordingly. For Key projection layers, we first compute head-wise CKA similarity, then reorder the heads to group the most similar ones together, and finally apply grouped SVD for compression. For Value projection layers, we apply standard SVD followed by offline calibration to adjust the low-rank factors. Then we fuse the right factor of the decomposition with the output projection matrix $W_{o}$ in the attention block. The modified matrices are then written back into the model. The final output is a modified model $\mathcal{M}^{\prime}$ that generates compressed KV caches.
@@ 4.1 Experimental Settings
Models. We evaluate our method on a range of widely adopted LLMs, including multiple generations of the LLaMA family: LLaMA-7B llama1 , LLaMA-2-7B llama2 , and LLaMA-2-13B-Chat llama2 . We also include instruction-tuned variants such as Mistral-7B-Instruct-v0.2 mistral and LongChat-7B-v1.5-32k li2023long to assess performance under extended context settings. These models cover both base and chat/instruction-following variants, ensuring a comprehensive evaluation across different architectures and use cases. Notably, Mistral-7B-Instruct-v0.2 mistral adopt Grouped Query Attention (GQA) ainslie2023gqa , while the others use standard Multi-Head Attention (MHA) vaswani2017attention .
Datasets and Evaluation. We assess the effectiveness of our method using both perplexity and task-specific accuracy. For language modeling evaluation, we report perplexity on WikiText2 wikitext2 , Penn Treebank (PTB)plotz2017benchmarking , and a subset of the C4 corpusraffel2020exploring . To evaluate reasoning and generalization capabilities, we measure zero-shot accuracy on six QA benchmarks: ARC-c, ARC-e clark2018think , Hellaswag zellers2019hellaswag , OBQA openbookqa , PIQA bisk2020piqa , and Winogrande winogrande . In addition, we adopt the LongBench benchmark LongBench to evaluate long-context performance, conducting experiments on eight diverse tasks that comprehensively test the model’s ability to handle extended contexts.
Baselines. We compare our method with Palu chang2024palu , a recent low-rank compression approach for KV cache. Specifically, we adopt its G-LRD variant for evaluation. To ensure a fair comparison, we follow the same group-wise decomposition setting and fix the group size to 4 in all experiments.
Implementation Details. All experiments are conducted using PyTorch paszke2019pytorch and Huggingface Transformers paszke1912imperative on a single NVIDIA A800 GPU with 80GB of memory. Following the setup in SVD-LLM svdllm , we apply a whitening transformation before performing SVD truncation. Specifically, we randomly select 256 samples from the WikiText-2 dataset as calibration data and use them both for whitening in the SVD step and for the offline calibration process in Value compression.
Table 1: Zero-shot performance comparison between ReCalKV and Palu chang2024palu under 50% to 70% compression ratios. Evaluation on three language modeling datasets (measured by perplexity ($\downarrow$)) and six zero-shot QA datasets (measured by both individual and average accuracy ($\uparrow$)).
@@
Mistral-7B-Instruct-v0.2 mistral
@@
Table 2: Evaluation results on LongBench LongBench , covering accuracy across 8 tasks and the overall average, comparing ReCalKV and Palu chang2024palu under 50%–70% KV cache compression ratios.
@@
LLaMA-2-13B-Chat llama2
@@
Mistral-7B-Instruct-v0.2 mistral
@@ 4.2 Main Results
Perplexity Results. We evaluate the language modeling capability of ReCalKV on three standard datasets: WikiText2 wikitext2 , PTB plotz2017benchmarking , and C4 c4 , using perplexity as the evaluation metric. As shown in Table 1, ReCalKV achieves lower perplexity than Palu chang2024palu on most compression ratios and model families. On LLaMA-2-7B llama2 , ReCalKV yields a perplexity of 5.83 on WikiText2 and 8.14 on C4 at 50% compression, compared to 6.02 and 8.72 from Palu, respectively. Similar trends are observed on LLaMA-7B llama1 and Mistral-7B mistral . Notably, on PTB, ReCalKV significantly outperforms Palu under aggressive compression. For instance, at 70% compression, the perplexity of Palu rises sharply to 211.33 on LLaMA-7B and 172.23 on LongChat-7B, while ReCalKV keeps it much lower at 71.47 and 109.38, respectively. These results demonstrate that ReCalKV maintains stronger language modeling ability under high compression. Even at 70% compression, perplexity degradation remains moderate, indicating better information preservation than existing low-rank methods.
Zero-shot Accuracy Results. In addition to perplexity, we evaluate ReCalKV on six zero-shot QA datasets, including OBQA openbookqa , HellaSwag hellaswag , PIQA bisk2020piqa , ARC-e arc , ARC-c arc , and Winogrande winogrande . Across all model families and compression levels, ReCalKV demonstrates strong resilience in accuracy. While both methods see a decline in performance as the compression ratio increases, Palu chang2024palu exhibits a significantly steeper drop. For instance, at 70% compression on LLaMA-2-7B llama2 , Palu’s average accuracy drops to 52.14%, whereas ReCalKV retains 59.90%. Similar robustness is observed on Mistral-7B mistral and LongChat-7B li2023long , where ReCalKV consistently delivers higher or comparable average accuracy under the same compression levels. These results highlight ReCalKV’s strong capability to preserve task performance even under aggressive KV cache size reductions. Moreover, the method’s consistent stability across a wide range of tasks further underscores its practical utility for efficient long-context inference in real-world applications.
Longbench Result. We further evaluate ReCalKV on LongBench LongBench , a benchmark designed to test long-context understanding across diverse tasks. As shown in Table 2, ReCalKV achieves higher average accuracy than Palu chang2024palu across nearly all model scales and compression ratios. The gap becomes especially pronounced at high compression levels (e.g., 70%), where Palu suffers significant degradation while ReCalKV maintains competitive performance. This demonstrates ReCalKV’s robustness and suitability for long-sequence inference under memory constraints.
@@ 4.3 Ablation study
To analyze the individual contributions of each component in ReCalKV, we conduct ablation studies on LLaMA-2-7B llama2 under a fixed 80% compression ratio. Table 3 reports perplexity on WikiText-2 wikitext2 , PTB plotz2017benchmarking , and C4 c4 , as well as accuracy on two downstream evaluation suites: the average accuracy over six zero-shot QA datasets (zero-shot Avg. Acc) and the average accuracy across eight tasks from the LongBench benchmark LongBench (LongBench Avg. Acc).
@@
Ablation on Offline Calibration. Comparing the first and third rows, we assess the effect of offline calibration alone. Perplexity on WikiText-2 improves from 9.34 to 8.91, while LongBench accuracy rises to 13.09%. This confirms that calibrating the SVD decomposition of the Value projection matrix using a small held-out dataset effectively improves the quality of approximation, thereby enhancing model performance and robustness across tasks.
@@ 4.4 Integrate with KV Cache Quantization
Since ReCalKV reduces the hidden dimension of the KV cache, it is inherently orthogonal and thus complementary to KV cache quantization techniques, which compress along the bitwidth axis.
@@
50% Palu chang2024palu 4 6.04 8.75
Palu chang2024palu 3 6.15 8.92
@@
60% Palu chang2024palu 4 6.84 10.77
Palu chang2024palu 3 7.01 11.06
@@
70% Palu chang2024palu 4 8.71 15.17
Palu chang2024palu 3 9.04 15.75
@@
This orthogonality enables seamless integration of the two methods for further compression. To assess their compatibility, we conduct experiments under various compression settings by combining ReCalKV with different quantization bitwidths (e.g., 4-bit and 3-bit), and varying the average rank to simulate different compression ratios. In our implementation, we adopt a per-token quantization scheme, allowing finer-grained control over dynamic range across sequence positions. Additionally, following Palu chang2024palu , we apply a randomized Hadamard transform prior to quantization to enhance robustness and reduce activation variance. Experimental results, as shown in Table 4, demonstrate that ReCalKV consistently outperforms Palu under the same compression ratio and bitwidth settings. For example, at a 60% compression ratio with 4-bit quantization, ReCalKV achieves a perplexity of 6.24 on WikiText-2, outperforming Palu’s 6.84. At 70% compression with 3-bit quantization, ReCalKV improves perplexity on C4 from 15.75 (Palu) to 10.41. These results confirm the effectiveness of combining ReCalKV with quantization for efficient KV cache compression.
5 Discussion
Limitations. While ReCalKV demonstrates high compression ratio with minimal accuracy drop, there remain several limitations to be addressed. First, despite our efforts to preserve performance, the compressed models still exhibit a certain degree of degradation in metrics such as perplexity and accuracy compared to the full KV cache baseline. This indicates that our method, like most lossy compression approaches, cannot fully maintain the original model behavior under high compression ratios. Second, ReCalKV currently focuses on hidden dimension reduction and does not incorporate token eviction strategies, which aim to reduce the KV cache length dimension. As these two directions are complementary, future work could explore the integration of ReCalKV with token eviction methods to further improve memory efficiency while maintaining performance.
Social Impacts. ReCalKV offers a practical and effective solution for compressing the KV cache in LLMs, significantly reducing memory consumption during long-context inference. This reduction in computational and memory requirements can lower both the deployment cost and energy consumption, thereby enabling wider access to LLMs in resource-constrained environments such as edge devices and developing regions. Although compression may lead to slight performance degradation, the trade-off is often acceptable in exchange for improved efficiency and scalability. By facilitating more sustainable and accessible AI deployment, ReCalKV contributes to the broader goal of inclusive and environmentally responsible AI development.

### v3

@@ 3.3 Offline Value Calibration
[…]
where $\mathbf{X}$ denotes the calibration dataset. Based on our analysis of Fisher Information, we observe that the Value projection matrix exhibits significantly higher Fisher Information compared to the Key projection matrix. This indicates that the Value projection matrix plays a more critical role in model performance. Therefore, we aim to minimize the approximation error $\mathcal{E}$ introduced during the compression of the Value projection to preserve model accuracy as much as possible. Inspired by Li et al. (2025), we observe that standard SVD decomposition does not always yield the lowest approximation error. To this end, we perform offline calibration of the decomposed matrices $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ using a small calibration dataset $\mathbf{X}$, aiming to further reduce the compression-induced approximation error. We first calibrate $\mathbf{L}_{v}$ by setting the derivative of the approximation error $\mathcal{E}$ with respect to $\mathbf{L}_{v}$ to zero:
@@
This yields an optimal closed-form solution for $\mathbf{R}_{v}$. Together, the calibrated $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$ form a refined low-rank approximation of the original Value projection matrix, effectively reducing compression-induced error without incurring any additional inference-time computation. The derivation, closed-form updates for $\mathbf{L}_{v}$ and $\mathbf{R}_{v}$, implementation details, related theoretical analysis, and the impact of different calibration datasets and dataset sizes are discussed in the supplementary material.
@@
[…]
@@
[…]
@@
[…]
@@ 3.4 ReCalKV Workflow
The overall pipeline of ReCalKV is outlined in Algorithm 1. Given a pre-trained model $\mathcal{M}$, calibration data $\mathcal{X}$, and a target compression ratio $\mathcal{R}$, ReCalKV applies differentiated compression strategies to Key and Value projection layers. We first compute layer-wise Fisher Information scores using the calibration data, following the strategy introduced in Palu (Chang et al., 2024), to estimate the relative importance of each layer and guide compression ratio allocation. Based on this, we allocate compression ratios to different layers accordingly. For Key projection layers, we first compute head-wise CKA similarity, then reorder the heads to group the most similar ones together, and finally apply grouped SVD for compression. For Value projection layers, we apply standard SVD followed by offline calibration to adjust the low-rank factors. Then we fuse the right factor of the decomposition with the output projection matrix $W_{o}$ in the attention block. The modified matrices are then written back into the model. The final output is a modified model $\mathcal{M}^{\prime}$ that generates compressed KV caches.
@@ 4.1 Experimental Settings
Models. We evaluate our method on a range of widely adopted LLMs, including multiple generations of the LLaMA family: LLaMA-7B (Touvron et al., 2023a), LLaMA-2-7B (Touvron et al., 2023b), and LLaMA-2-13B-Chat (Touvron et al., 2023b). We also include instruction-tuned variants such as Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) and LongChat-7B-v1.5-32k (Li et al., 2023) to assess performance under extended context settings. These models cover both base and chat/instruction-following variants, ensuring a comprehensive evaluation across different architectures and use cases. Notably, Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) adopt Grouped Query Attention (GQA) (Ainslie et al., 2023), while the others use standard Multi-Head Attention (MHA) (Vaswani, 2017). Additional results on the more recent LLaMA-3.1 model are provided in the supplementary file.
Datasets and Evaluation. We assess the effectiveness of our method using both perplexity and task-specific accuracy. For language modeling evaluation, we report perplexity on WikiText2 (Merity et al., 2017), Penn Treebank (PTB)(Plotz & Roth, 2017), and a subset of the C4 corpus(Raffel et al., 2020b). To evaluate reasoning and generalization capabilities, we measure zero-shot accuracy on six QA benchmarks: ARC-c, ARC-e (Clark et al., 2018b), Hellaswag (Zellers et al., 2019b), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), and Winogrande (Sakaguchi et al., 2020). In addition, we adopt the LongBench benchmark (Bai et al., 2023) to evaluate long-context performance, conducting experiments on eight diverse tasks that thoroughly evaluate long-context capability.
Baselines. We compare our method with Palu (Chang et al., 2024), a recent low-rank compression approach for KV cache. Specifically, we adopt its G-LRD variant for evaluation. For fair comparison, we adopt the same group-wise decomposition with a fixed group size of 4. Further comparisons with ASVD (Yuan et al., 2023) and EigenAttention (Saxena et al., 2024) are included in the supplementary.
Implementation Details. All experiments are conducted using PyTorch (Paszke et al., 2019b) and Huggingface Transformers (Paszke et al., 2019a) on a single NVIDIA A800 GPU with 80GB of memory. Following the setup in SVD-LLM (Wang et al., 2024), we apply a whitening transformation before performing SVD truncation. Specifically, we randomly select 256 samples from the WikiText-2 dataset as calibration data and use them both for whitening in the SVD step and for the offline calibration process in Value compression.
Table 1: Zero-shot performance comparison between ReCalKV and Palu (Chang et al., 2024) under 50% to 70% compression ratios. Evaluation on three language modeling datasets (measured by perplexity ($\downarrow$)) and six zero-shot QA datasets (measured by both individual and average accuracy ($\uparrow$)).
@@
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023)
@@
Table 2: Evaluation results on LongBench (Bai et al., 2023), covering accuracy across 8 tasks and the overall average, comparing ReCalKV and Palu (Chang et al., 2024) under 50%–70% KV cache compression ratios.
@@
LLaMA-2-13B-Chat (Touvron et al., 2023b)
@@
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023)
@@ 4.2 Main Results
Perplexity Results. We evaluate the language modeling capability of ReCalKV on three standard datasets—WikiText2 (Merity et al., 2017), PTB (Plotz & Roth, 2017), and C4 (Raffel et al., 2020a)—using perplexity as the metric. As shown in Table 1, ReCalKV achieves lower perplexity than Palu (Chang et al., 2024) on most compression ratios and model families. On LLaMA-2-7B (Touvron et al., 2023b), ReCalKV yields a perplexity of 5.83 on WikiText2 and 8.14 on C4 at 50% compression, compared to 6.02 and 8.72 from Palu, respectively. Similar trends are observed on LLaMA-7B (Touvron et al., 2023a) and Mistral-7B (Jiang et al., 2023). Notably, on PTB, ReCalKV significantly outperforms Palu under aggressive compression. For instance, at 70% compression, the perplexity of Palu rises sharply to 211.33 on LLaMA-7B and 172.23 on LongChat-7B, while ReCalKV keeps it much lower at 71.47 and 109.38, respectively. These results demonstrate that ReCalKV maintains strong language modeling ability under high compression. Even at 70% compression, perplexity remains moderate, suggesting better information retention than low-rank baselines.
Zero-shot Accuracy Results. In addition to perplexity, we evaluate ReCalKV on six zero-shot QA datasets, including OBQA (Mihaylov et al., 2018), HellaSwag (Zellers et al., 2019a), PIQA (Bisk et al., 2020), ARC-e (Clark et al., 2018a), ARC-c (Clark et al., 2018a), and Winogrande (Sakaguchi et al., 2020). Across all model families and compression levels, ReCalKV demonstrates strong resilience in accuracy. While both methods see a decline in performance as the compression ratio increases, Palu (Chang et al., 2024) exhibits a significantly steeper drop. For instance, at 70% compression on LLaMA-2-7B (Touvron et al., 2023b), Palu’s average accuracy drops to 52.14%, whereas ReCalKV retains 59.90%. Similar robustness is observed on Mistral-7B (Jiang et al., 2023) and LongChat-7B (Li et al., 2023), where ReCalKV consistently delivers higher or comparable average accuracy under the same compression levels. These results highlight ReCalKV’s strong capability to preserve task performance even under aggressive KV cache size reductions. Moreover, its stability across diverse tasks highlights its practicality for efficient long-context inference.
Longbench Result. We further evaluate ReCalKV on LongBench (Bai et al., 2023), a benchmark designed to test long-context understanding across diverse tasks. As shown in Table 2, ReCalKV achieves higher average accuracy than Palu (Chang et al., 2024) across nearly all model scales and compression ratios. The gap becomes especially pronounced at high compression levels (e.g., 70%), where Palu suffers significant degradation while ReCalKV maintains competitive performance across all benchmarks. This underscores ReCalKV’s robustness under memory constraints, enabling accurate and reliable long-context inference even at aggressive compression levels.
@@ 4.3 Ablation study
To analyze the individual contributions of each component in ReCalKV, we conduct ablation studies on LLaMA-2-7B (Touvron et al., 2023b) under a fixed 80% compression ratio. Table 3 reports perplexity on WikiText-2 (Merity et al., 2017), PTB (Plotz & Roth, 2017), and C4 (Raffel et al., 2020a), as well as accuracy on two downstream evaluation suites: the average accuracy over six zero-shot QA datasets (zero-shot Avg. Acc) and the average accuracy across eight tasks from the LongBench benchmark (Bai et al., 2023) (LongBench Avg. Acc).
@@
Ablation on Offline Value Calibration (OVC). Comparing the first and third rows, we assess the effect of offline calibration alone. Perplexity on WikiText-2 improves from 9.34 to 8.91, while LongBench accuracy rises to 13.09%. This confirms that calibrating the SVD decomposition of the Value projection matrix using a small held-out dataset effectively improves the quality of approximation, thereby enhancing model performance and robustness across tasks.
@@ 4.4 Integrate with KV Cache Quantization
50% Palu 4 6.04 8.75
Palu 3 6.15 8.92
@@
60% Palu 4 6.84 10.77
Palu 3 7.01 11.06
@@
70% Palu 4 8.71 15.17
Palu 3 9.04 15.75
@@
To evaluate the compatibility of ReCalKV with quantization, we combine it with 4-bit and 3-bit per-token quantization under varying average ranks to simulate different compression ratios. We also apply a randomized Hadamard transform before quantization, following Palu (Chang et al., 2024), to improve robustness. As shown in Table 4, ReCalKV consistently outperforms Palu under the same bitwidth and compression settings. For instance, at 60% compression with 4-bit quantization, ReCalKV achieves 6.24 perplexity on WikiText2 (vs. 6.84 for Palu); at 70% with 3-bit, it reduces C4 perplexity from 15.75 to 10.41. These results highlight the synergy between ReCalKV and quantization for efficient KV cache compression.
Figure 4: Latency speedup of ReCalKV relative to the baseline under various prompt lengths. Higher compression leads to greater acceleration, especially for longer prompts.
4.5 Inference Efficiency
To evaluate the practical runtime benefits of ReCalKV, we implement a custom fused attention kernel using Triton that integrates our low-rank compression for both Key and Value. For the Key path, we incorporate Head-wise Similarity-aware Reordering (HSR) as an online permutation step applied to each token during runtime. For the Value path, we perform offline matrix fusion to precompute and store a compact representation. The kernel supports rotary position embedding (RoPE) and is fully compatible with causal attention. We benchmark the latency of a single attention module on an NVIDIA A800 GPU across prompt lengths of 4K, 16K, and 65K, averaging over 100 runs per setting. As shown in Figure 4, ReCalKV achieves increasing latency speedups with higher KV compression and longer prompts, reaching up to 1.22×, 1.59×, and 1.80× improvements at 4K, 16K, and 65K respectively, under 70% compression. This trend confirms that our fused attention kernel becomes more effective as memory cost dominates, especially in long-context scenarios. These results validate the scalability and deployment efficiency of ReCalKV under strict memory budgets.